# DeferredGS: Decoupled and Editable Gaussian Splatting with Deferred Shading

### v1

@@ 4.8 Deferred Shading
In the decomposition and relighting process, we utilize the deferred shading technique instead of forward shading commonly used in previous methods. Here we ablate its influence on the relighting results in Fig. 8. Even if the geometry is enhanced by the normal field distillation, we can observe notable Gaussian shape-like blending artifacts when we apply the forward shading technique. By contrast, the deferred shading technique we use can successfully resolve this issue and produce smooth and realistic relighting results. We also evaluate it quantitatively in Table III and the relighting quality drops when it is not applied.

### v2

@@ 4.8 Deferred Shading
In the decomposition and relighting process, we utilize the deferred shading technique instead of forward shading commonly used in previous methods. Here we ablate its influence on the relighting results in Fig. 8. Even though the geometry is enhanced by the normal field distillation, we can still observe notable Gaussian shape—like blending artifacts when we apply the forward shading technique. By contrast, the deferred shading technique we use can successfully resolve this issue and produce smooth and realistic relighting results. We also evaluate it quantitatively in Table III and the relighting quality drops when it is not applied.